PLEASE I HAVE ANOTHER QUESTION. I BELIEVE THAT THIS ONE HAS A DISCRIMINANT
CALCULATE THE VALUE OF THE DISCRIMINANT OF X^2+4X+4=0
THIS PROBLEM IS FROM COLLEGE ALGEBRA WOOKBOOK

Answer by stanbon(75887)

CALCULATE THE VALUE OF THE DISCRIMINANT OF X^2+4X+4=0
-------------
b^2-4ac = 4^2-4*1*4 = 0
=========================
Cheers,
Stan H.